OpenStudy (anonymous):

help? What is true about the solutions of a quadratic equation when the radicand of the quadratic formula is zero? No real solutions Two identical rational solutions Two different rational solutions Two irrational solutions

OpenStudy (freckles):

\[ax^2+bx+c=0 \\ x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\] so when b^2-4ac=0 we have \[x=\frac{-b \pm \sqrt{0}}{2a}= \frac{-b \pm 0 }{2a}=\frac{-b}{2a}\]

OpenStudy (freckles):

so what do you think that means

OpenStudy (anonymous):

its the last one

OpenStudy (triciaal):

@zachleestone check again rational means can be expressed as a ratio of 2 integers irrational means not with integers

OpenStudy (anonymous):

ok.i see